# Gradient-free stochastic optimization for additive models

### v2

@@ 1 Introduction
In the present paper, we show that such a dimension reduction property fails to hold in the context of gradient-free optimization. We consider additive modeling in the problem of minimizing an unknown function $f:\mathbb{R}^{d}\to\mathbb{R}$ when only sequential evaluations of values of $f$ are available, corrupted with noise. We assume that $f$ is either strongly convex or satisfies the Polyak-Łojasiewicz (PL) condition [22, 18] and admits an additive representation as described above, where the components $f_{j}$ are $\beta$-Hölder.
@@
The setting that we consider belongs to the family of gradient-free (or zero-order) stochastic optimization problems, for which a rich literature is now available, see [17, 6, 19, 23, 15, 25, 10, 20, 4, 2, 5, 1, 8, 21] and the references therein. These papers did not assume any additive structure of $f$. It was proved in [23] that the minimax optimal rate of the optimization error, when $f$ is $\beta$-Hölder with $\beta\geq 2$ and satisfies the quadratic growth condition, is of the order $T^{-(\beta-1)/\beta}$ as function of the number of sequential queries $T$, to within an unspecified factor depending on the dimension $d$. Further developments were devoted to exploring the dependency of the minimax rate on $d$ assuming that $f$ is $\beta$-Hölder with $\beta\geq 2$ and is either strongly convex [15, 25, 4, 3, 1, 21] or satisfies the PL condition [1, 9]. In the PL case, unconstrained minimization was studied while the strongly convex case was analyzed both in constrained and unconstrained settings. A considerable progress was achieved though the complete solution valid for all $\beta\geq 2$ is not yet available. There exists a minimax lower bound for the class of $\beta$-Hölder and strongly convex functions, which scales as $dT^{-(\beta-1)/\beta}$ (proved in [25] for $\beta=2$ and Gaussian noise and in [2] for all $\beta\geq 2$ and more general noise; see also [1] for a yet more general lower bound). Moreover, the same rate is attained for $\beta=2$ under general conditions on the noise (no independence or zero-mean assumption), see [2]. Thus, for $\beta=2$ we know that the minimax rate is of the order $d\sqrt{T}$. For $\beta>2$, the literature provides different dependencies of the upper bounds on $d$ determined by the geometry of the $\beta$-Hölder condition. Thus, for the Hölder classes defined by pointwise Taylor approximation the best known upper bound is of the order $d^{2-1/\beta}T^{-(\beta-1)/\beta}$ when strongly convex functions are considered [3, 21]. On the other hand, for the Hölder classes defined by tensor-type conditions one can achieve the rate $d^{2-2/\beta}T^{-(\beta-1)/\beta}$ both for strongly convex functions and for PL functions [1]. Finally, the recent paper [31], assuming again strong convexity, deals with the class of functions that admit a Lipschitz Hessian. This represents a type of Hölder condition for $\beta=3$ and the paper [31] derives the upper rate $dT^{-2/3}$. We note that the lower bound of [2] with the rate $dT^{-(\beta-1)/\beta}$ is valid for all the above mentioned Hölder classes since this lower bound is obtained on additive functions that belong to all of them. Thus, the rate $dT^{-(\beta-1)/\beta}$ appears to be minimax optimal not only for $\beta=2$ but also for $\beta=3$ under a suitable definition of $3$-Hölder class of strongly convex functions.
@@
The main result of the present paper is to establish an upper bound with the rate $dT^{-(\beta-1)/\beta}$ for the optimization error in the noisy gradient-free setting over the class of $\beta$-Hölder functions satisfying the additive model and either the PL condition or the strong convexity condition. Together with the lower bound proved in [2], it implies that $dT^{-(\beta-1)/\beta}$ is the minimax optimal rate of the optimization error in this setting for all $\beta\geq 2$. This conclusion is quite surprising as it goes against the intuition acquired from the classical results on nonparametric estimation mentioned above. Indeed, it means that, at least for $\beta\in\{2,3\}$, there is no improvement in the rate neither in $T$ nor in $d$ when passing from the general $\beta$-Hölder model to the additive $\beta$-Hölder model. If any, an improvement can be obtained only in a factor independent of $T$ and $d$. Such a property may be explained by the fact that the optimization setting is easier than nonparametric function estimation in the sense that it aims at estimating a specific functional of the unknown $f$ (namely, its minimizer) rather than $f$ as a whole object. We also refer to a another somewhat similar fact in the gradient-free stochastic optimization setting. Specifically, there is no dramatic difference between the complexity of minimizing a strongly convex function with Lipschitz gradient, which corresponds to the case $\beta=2$ discussed above with the minimax rate $d\sqrt{T}$, and the complexity of minimizing a convex function with no additional properties, for which one can construct an algorithm converging with the rate between $d^{1.5}\sqrt{T}$ and $d^{1.75}\sqrt{T}$ (up to a logarithmic factor) as recently shown in [7].
@@ 2 Problem setup
Let $\Theta$ be a closed convex subset of $\mathbb{R}^{d}$. We consider the problem of minimizing an unknown function $f:\mathbb{R}^{d}\to\mathbb{R}$ over the set $\Theta$ based on noisy evaluations of $f$ at query points that can be chosen sequentially depending on the past observations. Specifically, we assume that at round $t\in\{1,\dots,T\}$ we can observe two noisy evaluations of $f$ at points $\boldsymbol{z}_{t},\boldsymbol{z}^{\prime}_{t}\in\mathbb{R}^{d}$, i.e.,
@@
where $\xi_{t},\xi^{\prime}_{t}$ are scalar noise variables and the query points $\boldsymbol{z}_{t},\boldsymbol{z}^{\prime}_{t}$ can be chosen depending on $\{\boldsymbol{z}_{i},\boldsymbol{z}^{\prime}_{i},y_{i},y^{\prime}_{i}\}_{i=1}^{t-1}$ and on a suitable randomization.
@@
Input: Constraint set $\Theta$, kernel function $K:[-1,1]\to\mathbb{R}$, step size $\eta_{t}>0$ and perturbation parameter $h_{t}>0$, for $t=1,\dots,T$
Initialization: Generate vectors $\boldsymbol{r}_{t}=(r_{t,1},\dots,r_{t,d})\in\mathbb{R}^{d}$ for $t=1,\dots,T$, where the components $r_{t,i}$ are independent and distributed uniformly from the interval $[-1,1]$, and choose $\boldsymbol{x}_{1}\in\Theta$
@@
       Observe $y_{t}=f(\boldsymbol{x}_{t}+h_{t}\boldsymbol{r}_{t})+\xi_{t}$ and $y^{\prime}=f(\boldsymbol{x}_{t}-h_{t}\boldsymbol{r}_{t})+\xi_{t}^{\prime}$
        // query
       for $j=1,\dots,d$ do
             $g_{t,j}=\frac{d}{2h_{t}}(y_{t}-y_{t}^{\prime})K(r_{t,j})$
@@
Algorithm 1 Zero-Order Stochastic Projected Gradient
In order to minimize $f$, we apply a version of projected gradient descent with estimated gradient presented in Algorithm 1. Let $\{\eta_{t}\}_{t=1}^{T}$ be a sequence of positive numbers and let $\{\boldsymbol{g}_{t}\}_{t=1}^{T}$ be a sequence of random vectors. Consider any fixed $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$ and let the vectors $\boldsymbol{x}_{t}$ for $t=2,\dots,T$ be defined by the recursion
@@
[…]
At each round $t$ of the algorithm, we generate a random vector $\boldsymbol{r}_{t}=(r_{t,1},\dots,r_{t,d})\in\mathbb{R}^{d}$, where the components $r_{t,j}$ are independent and distributed uniformly on $[-1,1]$. For $h_{t}>0$, we draw two noisy evaluations
@@ 3 Statement of the results
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-PL function. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of Algorithm 1 with $\Theta=\mathbb{R}^{d}$, and
@@
$\vspace{4mm}h_{t}$ $=\Big{(}\frac{3\bar{L}}{\alpha}\cdot\frac{\kappa\sigma^{2}}{L^{2}\kappa_{\beta}^{2}}\Big{)}^{\frac{1}{2\beta}}\cdot\begin{cases}t^{-\frac{1}{2\beta}}&\text{if $\eta_{t}=\frac{4}{\alpha t}$},\vspace{3mm}\\ T^{-\frac{1}{2\beta}}&\text{if $\eta_{t}=\frac{1}{18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}$}.\end{cases}$
@@
$\mathbf{E}\left[f(\boldsymbol{x}_{T})-f^{*}\right]\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}(f(\boldsymbol{x}_{1})-f^{*})+\frac{d}{\alpha}\left(\texttt{A}_{1}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{2}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right),$
where $\texttt{A}_{1},\texttt{A}_{2}>0$ depend only on $\beta$ and $\sigma^{2}$.
@@
Under the conditions of Theorem 1, if $T\geq(\bar{L}/\alpha)((\bar{L}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\alpha)^{\frac{\beta}{2}}/L)$, then
$\mathbf{E}\left[f(\boldsymbol{x}_{T})-f^{*}\right]\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}(f(\boldsymbol{x}_{1})-f^{*})+\texttt{A}\frac{d}{\alpha}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}},$
where $\texttt{A}>0$ depends only on $\beta$ and $\sigma^{2}$.
The bounds of Theorem 1 and Corollary 1 show how the rate of convergence depends on the parameters $T,d$ and $\alpha$ but also on the aspect ratio ${\bar{L}}/{\alpha}$. Moreover, in Corollary 1 the condition $T\geq Cd^{\beta/2}$, where $C>0$ is a constant, does not bring an additional restriction when $2\leq\beta\leq 3$ since the condition is weaker than giving the range of $T$, for which the bound makes sense. Indeed, for $2\leq\beta\leq 3$ the bound is greater than a constant independent of $T,d$ if $T\leq Cd^{\beta/2}$.
@@
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-strongly convex function on a compact convex subset $\Theta$ of $\mathbb{R}^{d}$ such that $\max_{\boldsymbol{x}\in\Theta}\left\lVert\nabla f(\boldsymbol{x})\right\rVert\leq G$. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of Algorithm 1 with
@@
$\mathbf{E}\left[f(\bar{\boldsymbol{x}}_{T})-f(\boldsymbol{x})\right]\leq\frac{18G^{2}d}{\alpha T}+\frac{d}{\alpha}\left(\texttt{A}_{1}L^{\frac{2}{\beta}}+\texttt{A}_{2}d\bar{L}^{2}(LT)^{-\frac{2}{\beta}}\right)T^{-\frac{\beta-1}{\beta}},$
where $\texttt{A}_{1},\texttt{A}_{2}>0$ depend only on $\beta$ and $\sigma^{2}$.
@@
Under the conditions of Theorem 2, if $T\geq(d\bar{L})^{\beta/2}L^{-2}$ then
@@
where $\texttt{A}>0$ does not depend on $T,d$ and $\alpha$.
@@
In view of strong convexity, Theorem 2 and Corollary 2 immediately imply the corresponding bounds on the estimation error $\mathbf{E}[\left\lVert\bar{\boldsymbol{x}}_{T}-\boldsymbol{x}^{*}\right\rVert^{2}]$, where $\boldsymbol{x}^{*}$ is the minimizer of $f$ on $\Theta$. Thus, under the assumptions of Corollary 2 we have
@@
    $H^{2}(P_{F_{t}(\cdot,\cdot)},P_{F_{t}(\cdot+v,\cdot+v)})\leq I_{0}v^{2}\,,\quad\quad|v|\leq v_{0}\enspace,$ (5)
    for some $0<I_{0}<\infty$, $0<v_{0}\leq\infty$. Here, $P_{F(\cdot,\cdot)}$ denotes the probability measure corresponding to the c.d.f. $F(\cdot,\cdot)$.

### v3

@@ 1 Introduction
In the present paper, we show that such a dimension reduction property fails to hold in the context of gradient-free optimization. We consider additive modeling in the problem of minimizing an unknown function $f:\mathbb{R}^{d}\to\mathbb{R}$ when only sequential evaluations of values of $f$ are available, corrupted with noise. We assume that $f$ is either strongly convex or satisfies the Polyak-Łojasiewicz (PL) condition [22, 18] and admits an additive representation as described above, where the components $f_{j}$ are $\beta$-Hölder. Functions satisfying the PL condition will be for brevity referred to as PL-functions.
@@
The setting that we consider belongs to the family of gradient-free (or zero-order) stochastic optimization problems, for which a rich literature is now available, see [17, 6, 19, 23, 15, 25, 10, 20, 4, 2, 5, 1, 8, 21] and the references therein. These papers did not assume any additive structure of $f$. It was proved in [23] that the minimax optimal rate of the optimization error, when $f$ is $\beta$-Hölder with $\beta\geq 2$ and satisfies the quadratic growth condition, is of the order $T^{-(\beta-1)/\beta}$ as function of the number of sequential queries $T$, to within an unspecified factor depending on the dimension $d$. Further developments were devoted to exploring the dependency of the minimax rate on $d$ assuming that $f$ is $\beta$-Hölder with $\beta\geq 2$ and is either strongly convex [15, 25, 4, 3, 1, 21] or satisfies the PL condition [1, 9]. In the PL case, unconstrained minimization was studied while the strongly convex case was analyzed both in constrained and unconstrained settings. A considerable progress was achieved though the complete solution valid for all $\beta\geq 2$ is not yet available. There exists a minimax lower bound for the class of $\beta$-Hölder and strongly convex functions, which scales as $dT^{-(\beta-1)/\beta}$ (proved in [25] for $\beta=2$ and Gaussian noise and in [2] for all $\beta\geq 2$ and more general noise; see also [1] for a yet more general lower bound). Moreover, there exists an algorithm attaining the same rate for $\beta=2$ under general conditions on the noise (no independence or zero-mean assumption), see [2]. Thus, for $\beta=2$ we know that the minimax rate is of the order $d/\sqrt{T}$. For $\beta>2$, the literature provides different dependencies of the upper bounds on $d$ determined by the geometry of the $\beta$-Hölder condition. Thus, for the Hölder classes defined by Taylor polynomial approximations the best known upper bound is of the order $d^{2-1/\beta}T^{-(\beta-1)/\beta}$ when strongly convex functions are considered [3, 21]. On the other hand, for the Hölder classes defined by tensor-type conditions one can achieve the rate $d^{2-2/\beta}T^{-(\beta-1)/\beta}$ both for strongly convex functions and for PL functions [1]. Finally, the recent paper [31], assuming again strong convexity, deals with the class of functions that admit a Lipschitz Hessian. This represents a type of Hölder condition for $\beta=3$ and the paper [31] derives the upper rate $dT^{-2/3}$. We note that the lower bound of [2] with the rate $dT^{-(\beta-1)/\beta}$ is valid for all the above mentioned Hölder classes since this lower bound is obtained on additive functions that belong to all of them. Thus, the rate $dT^{-(\beta-1)/\beta}$ appears to be minimax optimal not only for $\beta=2$ but also for $\beta=3$ under a suitable definition of $3$-Hölder class of strongly convex functions.
@@
The main result of the present paper is to establish an upper bound with the rate $dT^{-(\beta-1)/\beta}$ for the optimization error in the noisy gradient-free setting over the class of $\beta$-Hölder functions satisfying the additive model and either the PL condition or the strong convexity condition. Together with the lower bound proved in [2], it implies that $dT^{-(\beta-1)/\beta}$ is the minimax optimal rate of the optimization error in this setting for all $\beta\geq 2$. This conclusion is quite surprising as it goes against the intuition acquired from the classical results on nonparametric estimation mentioned above. Indeed, it means that, at least for $\beta\in\{2,3\}$, there is no improvement in the rate neither in $T$ nor in $d$ when passing from the general $\beta$-Hölder model to the additive $\beta$-Hölder model. If any, an improvement can be obtained only in a factor independent of $T$ and $d$. Such a property may be explained by the fact that the optimization setting is easier than nonparametric function estimation in the sense that it aims at estimating a specific functional of the unknown $f$ (namely, its minimizer) rather than $f$ as a whole object. We also refer to a another somewhat similar fact in the gradient-free stochastic optimization setting. Specifically, there is no dramatic difference between the complexity of minimizing a strongly convex function with Lipschitz gradient, which corresponds to the case $\beta=2$ discussed above with the minimax rate $d/\sqrt{T}$, and the complexity of minimizing a convex function with no additional properties, for which one can construct an algorithm converging with the rate between $d^{1.5}/\sqrt{T}$ and $d^{1.75}/\sqrt{T}$ (up to a logarithmic factor) as recently shown in [7].
@@ 2 Problem setup
Let $\Theta$ be a closed convex subset of $\mathbb{R}^{d}$. We consider the problem of minimizing an unknown function $f:\mathbb{R}^{d}\to\mathbb{R}$ over the set $\Theta$ based on noisy evaluations of $f$ at query points that can be chosen sequentially depending on the past observations. Specifically, we assume that at round $t\in\{1,\dots,T\}$ we observe two noisy evaluations of $f$ at points $\boldsymbol{z}_{t},\boldsymbol{z}^{\prime}_{t}\in\mathbb{R}^{d}$, i.e.,
@@
where $\xi_{t},\xi^{\prime}_{t}$ are scalar noise variables and the query points $\boldsymbol{z}_{t},\boldsymbol{z}^{\prime}_{t}$ can be chosen depending on $\{\boldsymbol{z}_{i},\boldsymbol{z}^{\prime}_{i},y_{i},y^{\prime}_{i}\}_{i=1}^{t-1}$ and on a possible randomization.
@@
Input: Constraint set $\Theta$, function $K:[-1,1]\to\mathbb{R}$, step size $\eta_{t}>0$, and perturbation parameter $h_{t}>0$, for $t=1,\dots,T$
Initialization: Generate vectors $\boldsymbol{r}_{t}=(r_{t,1},\dots,r_{t,d})\in\mathbb{R}^{d}$ for $t=1,\dots,T$, where the components $r_{t,i}$ are independent and distributed uniformly from the interval [1,1][-1,1], and choose $\boldsymbol{x}_{1}\in\Theta$
@@
   Obtain observations $y_{t}=f(\boldsymbol{x}_{t}+h_{t}\boldsymbol{r}_{t})+\xi_{t}$ and $y_{t}^{\prime}=f(\boldsymbol{x}_{t}-h_{t}\boldsymbol{r}_{t})+\xi_{t}^{\prime}$ for $j=1,\dots,d$ do
    $g_{t,j}:=\frac{d}{2h_{t}}(y_{t}-y_{t}^{\prime})K(r_{t,j})$
@@
Algorithm 1
In order to minimize $f$, we apply a version of projected gradient descent presented in Algorithm 1. Let $\{\eta_{t}\}_{t=1}^{T}$ be a sequence of positive numbers and let $\{\boldsymbol{g}_{t}\}_{t=1}^{T}$ be a sequence of random vectors. Consider any fixed $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$ and let the vectors $\boldsymbol{x}_{t}$ for $t=2,\dots,T$ be defined by the recursion
@@
$\int uK(u)du=1,~\int u^{j}K(u)du=0,\ j=0,2,3,\dots,\ell,~\text{and}~\kappa_{\beta}\equiv\int|u|^{\beta}|K(u)|du<\infty.$ (3)
Assume that $\kappa:=\int K^{2}(r)dr$ is finite. Functions $K$ satisfying these conditions are not hard to construct. In particular, a construction based on Legendre polynomials can be used, see, for example, [23, 29, 4].
At each round $t$ of the algorithm, we generate a random vector $\boldsymbol{r}_{t}=(r_{t,1},\dots,r_{t,d})\in\mathbb{R}^{d}$, where the components $r_{t,j}$ are independent and distributed uniformly on [1,1][-1,1]. For $h_{t}>0$, we draw two noisy evaluations
@@ 3 Statement of the results
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-PL function. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the realization of Algorithm 1 with $\Theta=\mathbb{R}^{d}$, and
@@
ht
[…]
$=\Big{(}\frac{3\bar{L}}{\alpha}\frac{\kappa\sigma^{2}}{{2}L^{2}\kappa_{\beta}^{2}}\Big{)}^{\frac{1}{2\beta}}\,\begin{cases}t^{-\frac{1}{2\beta}}&\text{if $\eta_{t}=\frac{4}{\alpha t}$},\vskip 8.53581pt\\ T^{-\frac{1}{2\beta}}&\text{if $\eta_{t}=\frac{1}{18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}$}.\end{cases}$
@@
$\mathbf{E}\left[f(\boldsymbol{x}_{T})-f^{*}\right]\leq{\frac{\texttt{A}_{0}}{T}(f(\boldsymbol{x}_{1})-f^{*})+\texttt{A}\frac{d}{\alpha}\left(\left(\frac{1}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{1}{\alpha T}\right)^{\frac{1}{\beta}}\right),}$
where $\texttt{A}_{0}=\max\big{(}\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}{\kappa}}{\alpha},1\big{)}$ and $\texttt{A}>0$ depends only on $L,\bar{L}$}, $\beta$, and $\sigma^{2}$.
@@
Under the conditions of Theorem 1, if $T\geq\alpha^{\beta/2-1}d^{\beta/2}$, then
$\mathbf{E}\left[f(\boldsymbol{x}_{T})-f^{*}\right]\leq\frac{\texttt{A}_{0}}{T}(f(\boldsymbol{x}_{1})-f^{*})+\texttt{A}\frac{d}{\alpha}\left(\frac{1}{\alpha T}\right)^{\frac{\beta-1}{\beta}},$
where $\texttt{A}>0$ is a constant that does not depend on $\alpha$, $d$, and $T$.
Note that in Corollary 1 the condition $T\geq Cd^{\beta/2}$, where $C>0$ is a constant, does not bring an additional restriction when $2\leq\beta\leq 3$ since the condition is weaker than giving the range of $T$, for which the bound makes sense. Indeed, for $2\leq\beta\leq 3$ the bound is greater than a constant independent of $T,d$ if $T\leq Cd^{\beta/2}$.
@@
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. Let $\Theta$ be a compact convex subset of $\mathbb{R}^{d}$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-strongly convex function such that $\max_{\boldsymbol{x}\in\Theta}\left\lVert\nabla f(\boldsymbol{x})\right\rVert\leq G$. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the realization of Algorithm 1 with
@@
$\mathbf{E}\left[f(\bar{\boldsymbol{x}}_{T})-f(\boldsymbol{x})\right]\leq\frac{36G^{2}d\kappa}{\alpha T}+\texttt{A}\frac{d}{\alpha}\left(1+dT^{-\frac{2}{\beta}}\right)T^{-\frac{\beta-1}{\beta}},$
where $\texttt{A}>0$ depends only on $L,\bar{L}$}, $\beta$, and $\sigma^{2}$.
@@
Under the conditions of Theorem 2, if $T\geq d^{\beta/2}$ then
@@
where $\texttt{A}>0$ is a constant that does not depend on $\alpha$, $d$, and $T$.
@@
In view of strong convexity, Theorem 2 and Corollary 2 immediately imply the corresponding bounds on the estimation error $\mathbf{E}[\left\lVert\bar{\boldsymbol{x}}_{T}-\boldsymbol{x}^{*}\right\rVert^{2}]$, where $\boldsymbol{x}^{*}$ is the minimizer of $f$ on $\Theta$ provided it is a global minimizer. Thus, under the assumptions of Corollary 2, if $\nabla f(\boldsymbol{x}^{*})=0$ we have
@@
    $H^{2}(P_{F_{t}(\cdot,\cdot)},P_{F_{t}(\cdot+v,\cdot+w)})\leq I_{0}\max(v^{2},w^{2})\,,\quad\quad|v|,|w|\leq v_{0},$ (5)
    for some $0<I_{0}<\infty$, $0<v_{0}\leq\infty$. Here, $P_{F(\cdot,\cdot)}$ denotes the probability measure corresponding to the cumulative distribution function $F(\cdot,\cdot)$.